Consider the following data

X 1 3 4 6 9 8 10

Y 1 8 15 33 75 70 95

The coefficient of correlation is

1. 0:976

2. 1:00

3. 0:82

4. 0:953

5. 0:88


1
Expert's answer
2021-09-06T14:44:14-0400

Solution:



X Values:

∑ = 41

Mean = 5.857

∑(X - Mx)2 = SSx = 66.857


Y Values:

∑ = 297

Mean = 42.429

∑(Y - My)2 = SSy = 8327.714


X and Y Combined:

N = 7

∑(X - Mx)(Y - My) = 728.429


R Calculation

r = ∑((X - My)(Y - Mx)) / "\sqrt{(SS_x)(SS_y)}"


r = 728.429 / "\sqrt{(66.857)(8327.714)}" = 0.9762

Thus, option 1 is correct.